• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kalyan

    Log entry time 11:19:07 on March 17, 2012

    Entry number 367122

    Followups:

    keyword=Update

    We lost both septum magnets around 9.30 am and at the same time we saw some He level issues with target. There were fire alarm and VESDA alarms. We called Ed and Jack - they are investigating the problem with septum magnet. From what I gathered so far, Ed thinks that we blew the coils on the septum (more later ..). The target magnet is ramped down and locked out to let people access the platform.



    A copy of this log entry has been emailed to: jpchen,slifer,camsonne,doug,brads,rom